OpenStudy (anonymous):

Determine whether the Mean Value Theorem can be applied to the function f(x) = lnx^2 on [1, sqr(e)]. If you can use the MVT, determine a value c that satisfies the theorem.

OpenStudy (anonymous):

mabe @just_one_last_goodbye or @magan @greenglasses @Alchemista @Vcarias123 @iamabarbiegirl @nincompoop @One098 @willreel @countrygirl1431 @BinkyBoo @samixboo @tom982 EmmaMink @quickstudent @quin100 @winters333

OpenStudy (freckles):

f(x)=ln(x^2) or f(x)=(ln(x))^2

OpenStudy (anonymous):

do you know what the mean value theorem is?

OpenStudy (anonymous):

Do you know what criteria have to be satisfied to be able to apply the MVT?

OpenStudy (anonymous):

f(x) = ln(x^2) And yes, I understand the (f(b) - f(a))/ (b - a), but that's what I'm having trouble with... would it be (1/ (sqrt(e) - 1)? What do I do from there?

OpenStudy (anonymous):

Before applying it, you need to see whether you can apply it. These need to be satisfied: 1. f is continuous on the closed interval [1, sqrt{e}] 2. f is differentiable on the open interval (1, sqrt{e})

OpenStudy (anonymous):

Ohh. I got yes for both. Is this the correct way to approach the problem? f'(x) = 2/x 2/x = 1/ (sqrt(e) - 1) c = 2(sqrt(e) -1)

OpenStudy (freckles):

looks good

OpenStudy (freckles):

now is c in the interval given?

OpenStudy (anonymous):

i believe so— it should be a positive number around 1.3 -sh

OpenStudy (freckles):

ok cool stuff